OpenStudy (anonymous):

if y varies inversely as x and y=12 when x=20 then what is y when x is 15

OpenStudy (texaschic101):

y = k/x (y = 12 and x = 20) 12 = k/20 -- multiply by 20 (20)(12) = k 240 = k y = 240/x (when x is 15) y = 240/15 y = 16